OpenStudy (anonymous):

Using the given points, determine Δy. (-3, -5) and (0, 10) a.Δy = 3 b.Δy = 5 c.Δy = 13 d.Δy = 15

OpenStudy (anonymous):

delta y = y2 - y1

OpenStudy (anonymous):

How much did y change? 10 - (-5)

OpenStudy (anonymous):

delta y means the change in Y, correct?

OpenStudy (pinksapphire):

What is 10-(-5)? If you have two negatives by each other you can just make them into positive like this: 10+5= So what's 10+5=?

OpenStudy (anonymous):

i tihnk 10+5=15

OpenStudy (anonymous):

@frebrezze what do u tink?

OpenStudy (anonymous):

he must have taken a break...

OpenStudy (pinksapphire):

10+5=15